What describes the lines represented by these equations?
X+y=9
X-y=9

Answer by TimothyLamb(4379)

Converting these lines to slope-intercept form: y = mx + b
---
L1: y = -x + 9
L2: y = x - 9
---
we can say that the lines are perpendicular, because their slopes are negative reciprocal: 1/1 and -1/1
---
Also, L1 and L2 intersect at (9, 0)
---
We can see this by substituting (x - 9) for y in L1 and solving for x (x=9), then substituting x=9 into either equation to find y=0.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php